A Complete Bibliography of Parallel and Distributed Computing Practices and Scalable Computing:

Practice and Experience

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

16 April 2022 Version 2.05

Title word cross-reference

0 [ICC02]. 1 [ICC02]. 3 [Mac00]. R [BSF98].

2000 [ZSnH01].

A-NETL [BYIO00]. Access [CPC00, HJM00]. Ad [ITT02, JVZ01]. Adaptation [JWP^{+}02]. Adapting [Key99]. Adaptive [BH99, KM02, KN02, Sch02, SM02a]. Aggressive [Agr99]. Air [GZ99]. algebra [Bar02]. Algorithm [CD00, HD02, ITT02, KSR02, BRPM98]. Algorithms [BF98, DFKL02, GZ99, GN98, GMMV00, PU02, Son00, TTKG02, Tro02, Bar02]. Analog [Akl02]. Analyser [MV00]. Analysis [ABD^{+}00, BGG^{+}02, HJHK02, HNW^{+}02, Kot00, WJT02]. Analyzing [GMMV00, TST02]. Annotations [GE98]. Anomaly [BTO^{+}01]. Ant [ITT02, KSR02]. Ant-colony [KSR02]. Antialiasing [MV00]. Application [AST02, Bes99, JWP^{+}02, QH00, SM02a, RJC98]. Applications [AKY^{+}02, BS99a, ES00, HB99, ITT02, Mac00, Mar02b, PK01, RLR00, SS00, TZVU02, WM01]. Applying [SRW01]. Approach [AKLL99, BSF98, GL99, KM02, eK98, SW00, SJC99]. Approaches [AJMJS02]. Archetypal [HIK99].

Architectural [Bad99, Key99]. Architecture [GFS99, SS00, CVLNP02]. Architectures [Gom98, HD02, NJN00, SW00, RJC98]. Array [HD02]. Arrays [SL01]. Assembly [Kut02]. Associative [KSSM00]. Asynchrony [SRU98]. ATM [NS98]. Atmospheric [Con99]. Automata [DC98, HLL^{+}98, MSC98]. Automatic [GE98, Yam98]. Availability [LLS01]. AVS [NH98].

Balancing [CD00, GL99, HB99, KN02, WJT02]. Band [Bes99]. Banded [ACDH99]. Based [AKY^{+}02, BTO^{+}01, KM02, LLS01, LS02, Omm01, SM02a, SM02b, TUO00, WZ99, dSBA99, vKdRST99, LG99, OOBG^{+}02, SP98, WCS99]. Behavior [TST02]. Benchmark [SWR^{+}00]. Benchmarks [ZSnH01]. benefit [AVC^{+}02]. Birge [HBL02]. Bit [SL01]. Bitstream [Kut02]. Block [PU02]. Block-Matching [PU02]. Board [Var99]. Body [AST02]. Book [Ang01, Ano98a, Ano98b, Ano99a, Ano99b, Ano99c, Ano00a, Ano00b, Ano01a, GH02, Kim01, Omm01, Pal01, Pan00a, Pan00b, Pan01]. Broadcast [vKdRST99]. Building [MW01]. Bus [LG99]. Bus-based [LG99].

C3 [BFG^{+}01]. Call [KL99]. Can [Gos01]. Capacities [NJN00]. Capacity [ABD^{+}00]. Capturing [SM02b]. CARPET [ST98]. Case [HNW^{+}02, HF98, PWPM99]. Cases [Gom98]. Caused [WJT02]. Cellular [DC98, MSC98]. Challenges [MW01, Mas02, Szy98]. Classifiers [LY01]. Cluster [BDDL01, CD00, FD99, HD02, KN02, BFG^{+}01]. Clusteral [KC00]. Clusters [Agr99, BS99b, Gos01, HPP02, WG99, dSBA99, Ano98d, MNT02, WJJ98]. collaborative [RMB^{+}98]. Colony [ITT02, KSR02]. Comeback [Red99]. Command [BFG^{+}01]. Communication [AL99, Agr99, Ano01b, CX01, GB99, Sch02, Ano98d, WJJ98]. Communications [CD00]. Comparing [GN98, ICC02]. Comparison [ACDH99]. Comparisons [HIK99]. Compensation [PU02]. Complex [MS01, MW01, SM02a, AVC^{+}02]. Complexity [BK00]. Component [KM02, PL00]. Component-Based [KM02]. Components [DdSL02, PWPM99]. Compound [WM01]. Compression [Kut02]. Computation [AL99, FNH99, FNS99, OBV02, NS98]. Computational [Akl02]. Computations [CPA00, HIK99, WZ99]. Computer [LY01, MP00]. Computers [Bes99, ST98, Tro02]. Computing [AKY^{+}02, AKLL99, AST02, Bad99, BS99b, DFKL02, FD99, Gos01, Kim01, LS02, MNT02, MLB99, MSC98, MB00, MW01, NJN00, PWPM99, QH00, Sch00, SRU98, SJC99, Szy98, PMM98]. Concepts [WLKB01]. Concerning [TST02]. Conciliating [Gir00]. Concurrent [Yam98, MMMS98]. Condition [HLL^{+}98]. Condition/Event [HLL^{+}98]. Conjugate [Bes99]. Connectionism [Gir00]. Connectivity [KSSM00]. Conservative [Ano01c, GLC^{+}01]. Considerations [HZLR02]. Consistency [SBM99]. Constructing [RLR00]. Consumption [JVZ01]. Control [BFG^{+}01, LLS01, Var99, Wel98, WH98, GFS99]. Cooperative [GL99]. CORBA [Ang01, ABD02, PWPM99]. Correlated [OSS01]. Critical [TW01]. Criticality [OSS01].

D [AKY^{+}02, Mac00]. Data [Ema99, HJM00]. Database [Pal01]. Databases [BDDL01]. Dataflow [SRU98]. DCOM [TUO00]. DCOM-Based [TUO00]. DDSCHED [BRPM98]. Deadline [HJHK02]. Debbuger [WCS99]. Debugger [KK01b]. Debugging [DLC01]. Decision [LY01]. Declustering [CK99]. decomposition [MMMS98]. Dedicated [MP00, Gos01]. Degradation [TST02]. delay [NS98]. Delayed [NsL98]. Delivery [PL00]. Demand [NX99]. dense [Bar02]. Derivation [GE98]. Descent [Akl01]. Design [BYIO00, BK00, HvR98, HPS00, NX99, PL00, Son00]. Designing [GH02, RMB^{+}98]. Detection [BTO^{+}01, CFMO01, LY01]. Development [Ano01c, GLC^{+}01, vKdRST99]. Diagonally [ACDH99]. DIET [CVLNP02]. Different [Kar01]. Differential [MV00]. Diffuse [RCJ00]. Diffusion [Ada00]. Digital [Gir00, MV00]. Dimensional [HD02]. Discrete [Akl01]. Distributed [AJMJS02, ABD02, Ano01b, ABD^{+}00, BL01, BDDL01, BK00, CFMO01, CPC00, CX01, DLC01, FNS99, GGCM99, GFS99, Gom98, HHH01, HJM00, HF98, KK01a, KK01b, LSDG98, Mac00, MS01, MMMM98, Mas02, MAT01, PZB02, QH00, Sch00, SW00, SJC99, Szy98, TUO00, TES99, Wel98, WH98, WZ99, WCS99, Yam98, Zal00, vKZ99, vKdRST99, BRPM98, PMM98, RMB^{+}98, RJC98]. distributed-object [RMB^{+}98]. Distribution [Var99]. Distributions [FNS99, Kar01]. DIWIDE [KK01b]. Do [Bol00, NH98]. Do-Loop [NH98]. DOGMA [SJC99]. Domain [Wel98]. Dominant [ACDH99]. Dynamic [SWR^{+}00, vKdRST99, BRPM98, RJC98]. DynBench [SWR^{+}00].

Earliest [HJHK02]. EDF [HZLR02]. Editor [Kow00, Tro01]. Editorial [Ano98c, Bad99, Bol00, Gos01, Key99, KL99, Mar02a, MW01, Mas02, Mil98, Pap01, Red99, Sch00, Son00, Szy98, Tro02, Wel98, Zal00]. Effect [KSSM00]. Effective [Yam98]. Efficent [TTKG02]. Efficient [BL01, CD00, CK99, KC00, NS98, Son00]. Electrocardiograph [AST02]. Electrode [AST02]. Embedded [TST02, Var99].

Empirical [QH00]. Encoding [Con99]. Energy [GFS99, JVZ01]. Engineering [WH98]. Ensuring [LLS01]. Enterprise [Kot00, NJN00]. Entry [SBM99]. Environment [Con99, LS02, ST98]. Environments [HNW^{+}02, Key99, KM02, WLKB01]. Estimate [Con99]. Estimating [NJN00]. Estimation [APT01]. Ethernet [HJHK02]. Evaluating [ZSnH01]. Evaluation [BYIO00, CPC00, CVLNP02]. Event [HLL^{+}98]. Events [WM01]. Evolution [WLKB01]. Excitable [Ada00]. Execution [KM01]. Expanding [LA02]. Experiences [DdSL02, FD99, FNS99, ITT02, MLB99, Mar02b]. Experimental [KSR02]. Experiments [GFS99]. Exploratory [Mac00]. Explorer [Nag00].

Facilitating [Kot00, PMM98].

Factorization [HBL02]. FARM [SM02a]. fast [Ano98d, WJJ98, TTKG02]. Faster [Tro02]. Fault [TES99]. Fault-Tolerant [TES99]. Feedback [SM02a]. Feedback-Based [SM02a]. FFTs [BSF98]. File [Akl02, CK99]. Filtering [SL01]. First [HJHK02, KP01]. Fixed [MV00]. Fixed-point [MV00]. Flow [STP02]. Fluid [STP02]. Formal [HvR98]. Fourier [TTKG02]. Frame [NsL98]. Framework [AVC^{+}02, KM01, KP01, RJC98, SP98]. Frequency [BH99]. Frequency-Adaptive [BH99]. Function [Con99]. Functional [BSF98, Bol00]. Functions [Var99, AVC^{+}02]. Furniture [Akl02]. Future [Bol00].

Gang [Kar01]. General [ACDH99]. Generalized [MSC98]. Generation [Var99]. Generic [MMMS98]. Geneva [DC98]. GP [TUO00]. Gracefull [TST02]. GRADE [PK01]. Gradient [Bes99]. Grammar [WZ99]. Graphics [Mac00]. GRID [Mar02a, AST02]. Group [GL99].

Handling [HJHK02]. Hardware [Gir00]. Heat [STP02]. Heterogeneous [AJMJS02, AKY^{+}02, AKLL99, DFKL02, KM02, KN02, LS02, MP00, QH00, RLR00, Bar02, MNT02]. Hierarchical [TES99, CVLNP02]. High [Bad99, BS99b, DFKL02, GFS99, WG99]. High-Energy [GFS99]. High-performance [Bad99]. Hilbert [Zin02]. HiPer [AKY^{+}02]. HiPer-D [AKY^{+}02]. HLA [MAT01]. HLA/RTI [MAT01]. Hoc [JVZ01, ITT02]. Hostile [Key99]. Hybrid [HLL^{+}98, RCJ00].

I/O [CK99]. If [Nag00]. II [TP00]. Image [Gos01, QH00, RLR00]. Imagery [GGCM99]. Implementation [BYIO00, BDDL01, HPP02, ITT02, NX99, TUO00, Ang01]. Implementations [Mar02b, ICC02]. implementing [RMB^{+}98]. Imprecise [FNH99]. Inclusion [Yam98]. Independent [HvR98]. Instruction [SS00]. Integrated [MMMM98]. Integrating [HLL^{+}98]. Integration [BF98, GN98, Sch02, Zin02]. Integrity [LLS01]. Inter [RCJ00]. Inter-reflection [RCJ00]. Interconnections [Kim01]. Interoperability [LM01]. Interoperable [KP01, WLKB01]. Interval [Sch02]. Interword [KSSM00]. Intra [WJT02]. Intra-Movie [WJT02]. Introduction [BS99b, CS00, DFKL02, FP00, FK00, KK01a, Kow00, eK98, LM01, Mar02b, PMM98, PZB02, TZVU02, Tro01, WH98, vKZ99]. Intrusion [BTO^{+}01, CFMO01, LY01]. inverse [MNT02]. Issue [BS99b, Kow00, Tro01, DFKL02, KK01a, LM01, Mar02b, PZB02, TZVU02]. Issues [AJMJS02, HHH01, MAT01, WJT02, dSBA99].

Java [Omm01, AJMJS02, KM02, MKT98, RMB^{+}98, RLR00, WCS99]. Java-based [WCS99]. Java-RMI [RMB^{+}98]. JAWS [HPS00]. Join [BH99]. Journal [Mil98, Ano98c].

Knapsack [ICC02]. KPROC [SS00].

Kronecker [TP99, TP00].

Laboratory [LLS01]. Landsat [GGCM99]. Landsat-TM [GGCM99]. Language [AKLL99, BYIO00, Yam98]. Languages [Bol00]. Large [ABD^{+}00, BK00, CPA00, GFS99, Kot00, LS99]. Large-Scale [Kot00]. Level [PL00, JWP^{+}02]. line [DLC01, PK01]. Linear [ACDH99, HBL02, OBV02, Bar02]. Load [CD00, GL99, HB99, KN02, WJT02]. Load-Balancing [CD00]. Logic [SRW01]. Loop [GE98, LS02, NH98]. LUX [MP00].

Mach [MKT98]. Machines

[Ano01b, BH99, CX01, Bar02]. Mainstream [Gos01]. Make [Red99]. Management [JWP^{+}02, MAT01, SM02a]. Manager [KW98, MLB99]. Managing [Mas02]. Mapping [AKY^{+}02, Ema99]. Massively [SL01]. Massively-Parallel [SL01]. Matching [PU02]. Matrices [OBV02, TP99, TP00]. Matrix [GZ99, MNT02]. Medical [STP02]. Memory [AL99, Ano01b, Ano01c, BL01, CX01, GLC^{+}01, eK98, LSDG98]. Mesh [HB99, KN02, KSR02, SL01]. Mesh-partitioning [KSR02]. message [Ano98d, WJJ98]. Messages [MMMM98]. Messaging [dSBA99]. Meta [PL00]. Meta-Level [PL00]. metacomputing [CVLNP02]. Method [CPC00, Yam98]. Methods [Bes99, Ema99, NS98]. MICO [Ang01]. Middleware [HJM00]. Migration [WG99]. Mission [TW01]. Mobile [ITT02, PL00]. Model [BGG^{+}02, FNH99, LA02, MB00, RMB^{+}98]. Modeling [OBV02]. Models [DFKL02, Ema99, GZ99, KC00]. Modular [HPP02]. Monitoring [DLC01, LM01, PK01, WLKB01]. Motion [PU02]. Move [Gos01]. Movie [WJT02]. Moving [Akl02]. MPEG [NsL98]. MPI [GGCM99, HD02, ICC02, WCS99, ZSnH01]. Multi [HD02, KM01]. Multi-Dimensional [HD02]. Multi-Execution [KM01]. Multicast [GB99, LS99]. Multicriteria [CPC00]. Multidimensional [BF98, DdSL02, GN98, OBV02]. Multilevel [SRW01]. Multimedia [Kut02]. Multiple [SBM99]. Multiple-Writer [SBM99]. Multiplication [TP99, TP00]. Multiprocessor [Ano01c, ES00, GLC^{+}01, SM02b, Ano98d, WJJ98]. Multiprocessors [HPP02]. Multithreaded [TTKG02]. Multithreading [SRU98]. Myrinet [OOBG^{+}02, dSBA99]. Myrinet-Based [dSBA99, OOBG^{+}02].

name [Pap01]. Narrow [ACDH99]. Narrow-Banded [ACDH99]. Naval [CFMO01]. NETL [BYIO00]. Network [BTO^{+}01, BS99a, IPM^{+}99, ITT02, OOBG^{+}02, RLR00, NS98, GH02]. Network-Based [BTO^{+}01]. Networks [AL99, ABD02, AKLL99, CK99, Con99, DFKL02, IPM^{+}99, JVZ01, LS99, LG99, MLB99]. Networks-of-Workstations [AL99]. Neural [Con99]. New-Generation [Var99]. NFS [OOBG^{+}02]. Non [Gos01]. Non-dedicated [Gos01]. Norms [Zin02]. Nothing [BH99]. NT [FD99]. Numeric [Tro02]. Numerical [STP02, Zin02]. Numerics [TZVU02].

O [CK99]. Object [BYIO00, CPC00, GE98, HJM00, HPS00, Sch00, RMB^{+}98]. Object-Oriented [BYIO00, GE98, HJM00, HPS00]. Objects [HF98]. Off [PWPM99]. Off-the-Shelf [PWPM99]. On-Board [Var99]. On-line [DLC01]. onto [AKY^{+}02]. Open [AJMJS02, Ang01, Wel98]. OpenMP [BGG^{+}02, HPP02, HD02, ICC02, ITT02, LA02, Mar02b, ZSnH01]. operational [MMMS98]. Optical [Kim01]. Optimal [AST02, Sch02]. Optimistic [OSS01]. Optimization [BK00, ITT02, Kot00, Kut02]. Optimizations [Agr99, CD00]. Optimize [CPC00]. Optimizing [Bar02]. Order [SL01]. Organized [OSS01]. Oriented [BYIO00, GE98, HJM00, HPS00]. Origin [ZSnH01]. Overhead [Ano01b, CX01]. Overlap [AL99]. Overlapped [PU02]. Overview [AJMJS02]. Ownership [MAT01].

P [PK01]. P-GRADE [PK01]. Padding [JVZ01]. Papers [Mil98]. Paradigm [Akl02, Bad99, eK98, vKdRST99]. Paradigms [HIK99, HD02]. Parallel [AJMJS02, ACDH99, BYIO00, Bes99, BF98, CPA00, CK99, Con99, DdSL02, DC98, Ema99, FD99, FNH99, GZ99, Gir00, GN98, GMMV00, Gos01, HBL02, HIK99, KK01a, Kim01, KN02, KC00, Kut02, MS01, MP00, MSC98, MB00, MV00, OBV02, Pal01, PZB02, PU02, Saj98, SS00, Sch02, SL01, SRU98, SJC99, Son00, ST98, SRW01, Szy98, TP99, TP00, TUO00, Tro02, TZVU02, WLKB01, WM01, WJT02, vKZ99, PMM98]. Part [TP00]. Partitioning [CPA00, HZLR02, SRW01, KSR02, SP98]. passing [Ano98d, WJJ98]. Path [GE98]. patterns [Saj98]. PCs [CD00]. PDE [HIK99]. Performance [ABD^{+}00, BGG^{+}02, BK00, BS99b, DFKL02, HHH01, HPS00, Kar01, KM01, LS02, LSDG98, MS01, Nag00, NH98, OOBG^{+}02, PL00, SW00, WG99, WM01, ZSnH01, Bad99, SP98]. performance-based [SP98]. PFDDAA [MV00]. Physics [GFS99]. Pipelining [BL01]. Planning [ABD^{+}00]. Platform [FD99, HvR98]. Platform-Independent [HvR98]. Point [Con99, MV00]. Pollution [GZ99]. Portability [MB00]. Positioning [AST02]. Preconditioned [Bes99]. Preconditioners [Bes99]. Prediction [LS02]. Preface [DK99]. Prefix [SS00]. Primes [FNS99]. Principle [WZ99]. Problem [ICC02, KSR02, ZSnH01]. Problems [ST98, Wel98]. Process [ABD02]. Processes [HNW^{+}02]. Processing [ES00, GGCM99, Gos01, KSSM00, QH00, RLR00]. Processor [SL01]. Processors [Ada00]. Product [TP99, TP00]. program [SP98]. Programming [BS99a, Bol00, Ema99, HIK99, Mac00, ST98, Saj98, Omm01]. Programs [BGG^{+}02, GE98, HBL02, KK01a]. Projection [Ema99]. Properties [WM01]. Protocol [TES99]. Protocols [Ano01c, GLC^{+}01, HJHK02]. Provision [Gos01]. PVM [WCS99].

Qi [HBL02]. QOS [JWP^{+}02]. QOS-level [JWP^{+}02]. Quadrature [DdSL02]. Quality [KK01a]. Quantized [HZLR02].

Radio [vKdRST99]. Radiosity [CPA00]. Radix [BSF98]. Radix- [BSF98]. RAM [IPM^{+}99]. Rank [SL01]. Ray [GMMV00, MP00, RCJ00]. Reaction [Ada00]. Reaction-Diffusion [Ada00]. Real [Akl01, APT01, BTO^{+}01, BS99a, Bol00, CFMO01, ES00, FNH99, GB99, GFS99, Gom98, GE98, HJHK02, HvR98, HNW^{+}02, JWP^{+}02, KM02, MLB99, MMMM98, MW01, MKT98, NsL98, PZB02, PWPM99, SWR^{+}00, SM02a, SM02b, TW01, TES99, Var99, WZ99, Zal00, dKAE98, vKZ99, vKdRST99, AVC^{+}02, BRPM98, NS98]. Real- [GE98]. Real-Time [APT01, BTO^{+}01, BS99a, CFMO01, ES00, FNH99, GB99, GFS99, HJHK02, HvR98, HNW^{+}02, JWP^{+}02, KM02, MMMM98, MKT98, NsL98, PZB02, PWPM99, SWR^{+}00, SM02a, SM02b, TW01, TES99, Var99, Zal00, dKAE98, vKZ99, vKdRST99, Gom98, MW01, AVC^{+}02, BRPM98, NS98]. Realistic [ZSnH01]. Referees [Pap00].

Refinement [KN02]. Reflection [MMMS98, RCJ00]. Reliability [SM02b]. Reliable [IPM^{+}99]. Remote [HJM00]. Remotely [GGCM99]. Rendering [CD00]. Resource [CPC00, KW98, MLB99, SM02a]. Response [CFMO01]. Review [Ang01, GH02, Kim01, Omm01, Pal01]. Reviews [Ano98a, Ano98b, Ano99a, Ano99b, Ano99c, Ano00a, Ano00b, Ano01a, Pan00a, Pan00b, Pan01]. Ring [TES99]. RMI [RMB^{+}98]. Rough [WZ99]. RTI [MAT01]. Run [KP01]. Run-time [KP01].

Sampling [RCJ00]. Scalability [Nag00, OOBG^{+}02, SW00]. Scalable [ES00, MB00, Son00]. Scalar [APT01]. Scale [BK00, Kot00]. Scenario [Nag00]. Schedulability [SM02b]. Schedulability-Reliability [SM02b]. Scheduler [FNH99, SM02b]. Scheduling [CPA00, HZLR02, HJHK02, Kar01, LG99, LS02, MMMM98, QH00, RCJ00, dKAE98, BRPM98]. Schemes [JVZ01, NsL98, STP02]. Schmidt [Zin02]. Scientific [Bol00, MLB99, ST98]. Secure [Akl02]. Security [TW01]. Self [OSS01]. Self-Organized [OSS01]. Semantics [MMMS98]. Semi [PK01]. Semi-on-line [PK01]. SEMPA [MLB99]. Sense [Ada00]. Sensed [GGCM99]. Sensor [SM02a]. Server [MKT98, NJN00, Omm01]. Servers [NX99, WJT02]. Service [LLS01]. SGI [ZSnH01]. Shared [AL99, Ano01c, BL01, BH99, GLC^{+}01, KW98, eK98, LSDG98]. Shared-Memory [AL99]. Sharing [HNW^{+}02]. Shelf [PWPM99]. Signal [ES00]. Signals [OBV02]. SIMA [RLR00]. SIMD [Red99]. Simuation [HJHK02]. Simulation [DC98, MAT01, OSS01, SRW01]. Simulations [Nag00, STP02]. Single [Gos01, SL01]. Single-Bit [SL01]. Size [Kar01]. Sizes [ZSnH01]. Skewness [WJT02]. SMP [HD02]. SMPs [WG99].

Soft [KM02]. Software [BL01, Gom98, KP01, SW00, Zal00, dSBA99, vKdRST99, RJC98]. Solvers [ACDH99]. Solving [ST98]. Some [DdSL02, Mil98, MAT01]. Source [Ang01]. Sparse [GZ99]. SPE [SW00]. Special [BS99b, Kow00, Tro01, DFKL02, KK01a, LM01, Mar02b, PZB02, TZVU02]. Specifying [WM01]. Spread [Con99]. stage [HBL02]. Statically [AKY^{+}02]. Steepest [Akl01]. Stochastic [HBL02]. Strategies [Ema99, QH00]. Structured [OBV02]. Study [KSR02, QH00]. Suite [BFG^{+}01, SWR^{+}00]. Super [APT01]. Super-Scalar [APT01]. Supercomputing [KL99]. Superstep [Ano01c, GLC^{+}01]. Support [BL01, Zal00]. Supporting [DLC01, WCS99]. Surface [AST02]. Survey [dSBA99]. Switched [HJHK02]. Synchronization [LA02]. Synchronizer [BDDL01]. System [AKY^{+}02, APT01, GFS99, Gos01, QH00, Ano98d, RMB^{+}98, WJJ98]. Systems [AJMJS02, Ano01c, ACDH99, ABD^{+}00, BL01, BK00, CFMO01, CPC00, GLC^{+}01, HHH01, HvR98, HLL^{+}98, JWP^{+}02, KK01a, Kot00, LM01, LSDG98, MS01, MMMM98, MW01, Mas02, NX99, OSS01, PZB02, SWR^{+}00, SM02a, SM02b, TW01, TST02, TES99, Var99, Wel98, WH98, WLKB01, Yam98, dKAE98, dSBA99, vKZ99, vKdRST99, AVC^{+}02, MMMS98]. Systolic [SS00].

Task [LG99, QH00]. Tasks [MMMM98]. Techniques [AKY^{+}02, Pal01]. Technology [MS01]. Thoughts [Mil98]. Thread [WG99]. Three [HBL02, HIK99]. Three-stage [HBL02]. Time [Akl01, APT01, BTO^{+}01, BS99a, BDDL01, CFMO01, ES00, FNH99, GB99, GFS99, GE98, HJHK02, HvR98, HNW^{+}02, JWP^{+}02, KM02, MMMM98, MKT98, NsL98, PZB02, PWPM99, SWR^{+}00, SM02a, SM02b, TW01, TES99, Var99, WZ99, Zal00, dKAE98, vKZ99, vKdRST99, AVC^{+}02, BRPM98, Gom98, KP01, MW01, NS98]. Time-Sharing [HNW^{+}02]. Timed [HLL^{+}98, Yam98]. TM [GGCM99]. Tolerant [TES99]. Tool [BFG^{+}01, LM01, RLR00, WLKB01, Zal00]. Tools [DFKL02, KP01]. Tracer [RCJ00]. Tracing [GMMV00, MP00]. Tracking [HD02]. Tradeoff [SM02b]. Traffic [DC98, JVZ01, NS98]. Trafic [HJHK02]. Transaction [ABD^{+}00]. Transfer [Akl02, STP02]. Transform [TTKG02]. Transmission [NsL98]. Transposition [HD02]. Tree [LY01]. Trends [Sch00]. triangular [MNT02]. Tuning [KM01]. Tuplespace [KW98]. Twin [FNS99].

Unconventional [Ada00]. Unstructured [HB99]. Use [Gom98]. Using [DC98, GGCM99, IPM^{+}99, Kim01, NH98, WM01, Yam98, AVC^{+}02, MAT01, RMB^{+}98, Saj98]. Utilization [AKY^{+}02]. Utilization-Based [AKY^{+}02].

Vacancy [HD02]. Value [SM02b]. Value-Based [SM02b]. Vector [TP99, TP00]. Verification [HLL^{+}98, Yam98, vKdRST99]. Video [NX99, PL00, WJT02, dKAE98]. Video-on-Demand [NX99]. Videos [NsL98]. Virtual [BDDL01, HF98, KW98, eK98]. visual [Saj98]. Visualization [KC00, LSDG98, NH98]. Volume [CD00, KC00, Pap00].

Wakeup [KL99]. Way [HvR98]. WCET [APT01]. WDM [LS99]. Web [LLS01, SM02a]. Web-Based [LLS01]. What-If [Nag00]. Will [Red99]. Windows [FD99]. Wireless [GB99, GH02, JVZ01]. within [KM02, RCJ00]. Work [GL99]. workstation [Ano98d, WJJ98]. Workstations [AL99, Agr99, BDDL01, CK99, IPM^{+}99, LG99, MLB99, OOBG^{+}02]. World [MLB99]. Worst [HNW^{+}02]. Worst-Case [HNW^{+}02]. Writer [SBM99]. Writing [Mil98].

References

Aries:2000:CPP

[ABD^{+}00] James Aries, Marc Brittan, Eric Dillon, Alan Korncoff, Janusz Kowalik, and John Lixvar. Capacity planning for and performance analysis of large distributed transaction systems. Parallel and Distributed Computing Practices, 3(4):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol03/vol03no4abs.html#aries.

Amar:2002:TDP

[ABD02] Abdelkader Amar, Pierre Boulet, and Jean-Luc Dekeyser. Towards distributed process networks with CORBA. Parallel and Distributed Computing Practices, 5(4):??, December 2002. CODEN ???? ISSN 1097-2803.

Arbenz:1999:CPS

[ACDH99] P. Arbenz, A. Cleary, J. Dongarra, and M. Hegland. A comparison of parallel solvers for diagonally dominant and general narrow-banded linear systems. Parallel and Distributed Computing Practices, 2(4):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol02/vol02no4abs.html#arbenz.

Adamatzky:2000:RDE

[Ada00]  A. Adamatzky.  Reaction-

diffusion and excitable processors: A sense of the unconventional. Parallel and Distributed Computing Practices, 3(2):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol03/vol03no2abs.html#adamatzky.

Agrawal:1999:ACO

[Agr99] G. Agrawal. Aggressive communication optimizations for clusters of workstations. Parallel and Distributed Computing Practices, 2(2):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol02/vol02no2abs.html#gagan.

Al-Jaroodi:2002:OPD

[AJMJS02] Jameela Al-Jaroodi, Nader Mohamed, Hong Jiang, and David R. Swanson. An overview of parallel and distributed Java for heterogeneous systems: Approaches and open issues. Parallel and Distributed Computing Practices, 5(4):??, December 2002. CODEN ???? ISSN 1097-2803.

Akl:2001:DSD

[Akl01] Selim G. Akl. Discrete steepest descent in real time. Parallel and Distributed Computing Practices, 4(3):??, September 2001. CODEN ???? ISSN 1097-2803.

Akl:2002:SFT

[Akl02] Selim G. Akl. Secure file transfer: A computational analog to the furniture moving paradigm. Parallel and Distributed Computing Practices, 5(2):193–203, June 2002. CODEN ???? ISSN 1097-2803.

Arapov:1999:LAC

[AKLL99] D. Arapov, A. Kalinov, A. Lastovetsky, and I. Ledovskih. A language approach to computing on heterogeneous networks. Parallel and Distributed Computing Practices, 2(3):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol02/vol02no3abs.html#arapov.

Ali:2002:UBT

[AKY^{+}02] Shoukat Ali, Jong-Kook Kim, Yang Yu, Shriram B. Gundala, Sethavidh Gertphol, Howard Jay Siegel, Anthony A. Maciejewski, and Viktor Prasanna. Utilization-based techniques for statically mapping heterogeneous applications onto the HiPer-D heterogeneous computing system. Parallel and Distributed Computing Practices, 5 (4):??, December 2002. CODEN ???? ISSN 1097-2803.

Abdelrahman:1999:OCC

[AL99] T. S. Abdelrahman and G. Liu.

Overlap of computation and communication on shared-memory networks-of-workstations. Parallel and Distributed Computing Practices, 2(2):??, ????

1999. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol02/vol02no2abs.html#tarek.

Angryk:2001:BRM

[Ang01]  Rafal Angryk.  Book re-

view: MICO, An Open Source CORBA Implementation. Parallel and Distributed Computing Practices, 4(3):??, September 2001. CODEN ???? ISSN 1097-2803.

Anonymous:1998:BRa

[Ano98a]  Anonymous.  Book reviews.

Parallel and Distributed Computing Practices, 1(3):??, ???? 1998. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol01/vol01no3reviews.html.

Anonymous:1998:BRb

[Ano98b]  Anonymous.  Book reviews.

Parallel and Distributed Computing Practices, 1(4):??, ???? 1998. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol01/vol01no4reviews.html.

Anonymous:1998:EWN

[Ano98c] Anonymous. Editorial: Why a new journal? Parallel and Distributed Computing Practices, 1(1):??, ???? 1998. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol01/vol01no1editorial.html.

Anonymous:1998:NFM

[Ano98d] Anonymous. A new fast message passing communication system for multiprocessor workstation clusters. Parallel and Distributed Computing Practices, 1(4):??, ???? 1998. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol01/vol01no4abs.html#zhen.

Anonymous:1999:BRa

[Ano99a]  Anonymous.  Book reviews.

Parallel and Distributed Computing Practices, 2(1):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol02/vol02no1reviews.html.

Anonymous:1999:BRb

[Ano99b]  Anonymous.  Book reviews.

Parallel and Distributed Computing Practices, 2(2):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol02/vol02no2reviews.html.

Anonymous:1999:BRc

[Ano99c]  Anonymous.  Book reviews.

Parallel and Distributed Computing Practices, 2(4):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol02/vol02no4reviews.html.

Anonymous:2000:BRa

[Ano00a]  Anonymous.  Book reviews.

Parallel and Distributed Computing Practices, 3(1):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol03/vol03no1reviews.html.

Anonymous:2000:BRb

[Ano00b]  Anonymous.  Book reviews.

Parallel and Distributed Computing Practices, 3(2):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol03/vol03no2reviews.html.

Anonymous:2001:BR

[Ano01a]  Anonymous.  Book reviews.

Parallel and Distributed Computing Practices, 4(1):??, ???? 2001. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol04/vol04no1reviews.html.

Anonymous:2001:COD

[Ano01b] Anonymous. Communication overhead on distributed memory machines. Parallel and Distributed Computing Practices, 4(1):??, ???? 2001. CODEN ???? ISSN 1097-

2803.  URL http://www.
cs.okstate.edu/~pdcp/vols/
vol04/vol04no1abs.html#chen.

Anonymous:2001:DCS

[Ano01c] Anonymous. The development of conservative superstep protocols for shared memory multiprocessor systems. Parallel and Distributed Computing Practices, 4(1):??, ???? 2001. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol04/vol04no1abs.html#gan.

Aljifri:2001:EWS

[APT01] Hassan Aljifri, Alexander Pons, and Moiez Tapia. The estimation of the WCET in superscalar real-time system. Parallel and Distributed Computing Practices, 4(2):??, ???? 2001. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol04/vol04no2abs.html#aljifri.

Avbelj:2002:GCA

[AST02] V. Avbelj, M. Sterk, and Roman Trobec. A Grid computing application: Optimal electrode positioning in body surface electrocardiograph. Parallel and Distributed Computing Practices, 5 (3):??, September 2002. CO-DEN ???? ISSN 1097-2803.

Andrews:2002:FUB

[AVC^{+}02] David Andrews, Ravi Vemuri, David M. Chelberg, David Fleeman, David Parrott, Lonnie R. Welch, and Scott Brandt. A framework for using benefit functions in complex real-time systems. Parallel and Distributed Computing Practices, 5 (1):41–50, March 2002. CODEN ???? ISSN 1097-2803.

Bader:1999:ENA

[Bad99] David A. Bader. Editorial:

A new, architectural paradigm for high-performance computing. Parallel and Distributed Computing Practices, 2(2):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol02/vol02no2editorial.html.

Barbosa:2002:ODL

[Bar02]  Jorge Barbosa.  Optimizing

dense linear algebra algorithms on heterogeneous machines. Parallel and Distributed Computing Practices, 5(4):??, December 2002. CODEN ???? ISSN 1097-2803.

Boukerche:2001:IVT

[BDDL01] A. Boukerche, S. K. Das, A. Datta, and T. E. LeMaster. Implementation of a virtual time synchronizer for distributed databases on a cluster of workstations. Parallel and Distributed Computing Practices, 4(1):??, ???? 2001. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol04/vol04no1abs.html#boukerche.

Besson:1999:BPA

[Bes99] O. Besson. Band preconditioners: Application to preconditioned conjugate gradient methods on parallel computers. Parallel and Distributed Computing Practices, 2(4):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol02/vol02no4abs.html#besson.

Bull:1998:PAM

[BF98] J. M. Bull and T. L. Freeman. Parallel algorithms for multidimensional integration. Parallel and Distributed Computing Practices, 1(1):??, ????

1998. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol01/vol01no1abs.html#bull.

Brim:2001:CCC

[BFG^{+}01] Michael Brim, Ray Flanery, Al Geist, Brian Luethke, and Stephen L. Scott. Cluster Command and Control (C3) tool suite. Parallel and Distributed Computing Practices, 4(4):??, December 2001. CODEN ???? ISSN 1097-2803.

Blanco:2002:PMA

[BGG^{+}02] V. Blanco, L. García, J. A.

González, C. Rodríguez, and G. Rodríguez. A performance model for the analysis of OpenMP programs. Parallel and Distributed Computing Practices, 5(2):139–151, June 2002. CODEN ???? ISSN 1097-2803.

Bamha:1999:FAJ

[BH99] M. Bamha and G. Hains.

Frequency-adaptive join for shared nothing machines. Parallel and Distributed Computing Practices, 2(3):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol02/vol02no3abs.html#bamha.

Brittan:2000:POC

[BK00] Marc Brittan and Janusz Kowalik. Performance, optimization and complexity in the design of large scale distributed systems. Parallel and Distributed Computing Practices, 3(4):??, ????

2000. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol03/vol03no4abs.html#brittan.

Balasubramanian:2001:ESP

[BL01] Karthik Balasubramanian and David K. Lowenthal. Efficient support for pipelining in software distributed shared memory systems. Parallel and Distributed Computing Practices, 4(2):??, ???? 2001. CO-DEN ???? ISSN 1097-

                                        2803.  URL http://www.
                                        cs.okstate.edu/~pdcp/vols/
                                        vol04/vol04no2abs.html#balasubramanian.

Bollman:2000:EDF

[Bol00] D. Bollman. Editorial: Do functional languages have a real future in scientific programming? Parallel and Distributed Computing Practices, 3(4):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol03/vol03no4editorial.html.

Bhattacharjee:1998:DDD

[BRPM98] Anup K. Bhattacharjee, K. Ravindranath, A. Pal, and R. Mall. DDSCHED: A distributed dynamic real-time scheduling algorithm. Parallel and Distributed Computing Practices, 1(4):??, ???? 1998. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol01/vol01no4abs.html#mal.

Benzekri:1999:RTN

[BS99a] A. Benzekri and P. Sarafoglu. Real-time network applications programming. Parallel and Distributed Computing Practices, 2(1):??, ???? 1999. CO-DEN ???? ISSN 1097-

2803.  URL http://www.
cs.okstate.edu/~pdcp/vols/
vol02/vol02no1abs.html#benzekri.

Buyya:1999:HPC

[BS99b] Rajkumar Buyya and Clemens Szyperski. High performance computing on clusters?: Introduction to the special issue. Parallel and Distributed Computing Practices, 2(2):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol02/vol02no2introduction.
html.

Bollman:1998:FAR

[BSF98] D. Bollman, J. Seguel, and J. Feo. A functional approach to radix-R FFTs. Parallel and Distributed Computing Practices, 1(1):??, ???? 1998. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol01/vol01no1abs.html#bollman.

Balupari:2001:RTN

[BTO^{+}01] Ravindra Balupari, Brett Tjaden, Shawn Ostermann, Marina Bykova, and Aaron Mitchell. Real-time network-based anomaly intrusion detection. Parallel and Distributed Computing Practices, 4(2):??, ???? 2001. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol04/vol04no2abs.html#balupari.

Baba:2000:DIE

[BYIO00] T. Baba, T. Yoshinaga, Y. Iwamoto, and K. Ootsu. Design, implementation and evaluation of a parallel object-oriented language A-NETL. Parallel and Distributed Computing Practices, 3(2):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol03/vol03no2abs.html#baba.

Chaussumier:2000:COE

[CD00] F. Chaussumier and F. Desprez. Communications optimizations and efficient loadbalancing for a volume rendering algorithm on a cluster of PCs. Parallel and Distributed Computing Practices, 3(3):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol03/vol03no3abs.html#chaussumier.

Chappell:2001:IDR

[CFMO01] B. L. Chappell, G. A. Fink, D. T. Marlow, and K. F. O’Donoghue. Intrusion detection and response for realtime distributed naval systems. Parallel and Distributed Computing Practices, 4(2):??, ???? 2001. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol04/vol04no2abs.html#chappell.

Cheng:1999:FDE

[CK99] H-.Y. Cheng and C-.T. King. File declustering for efficient parallel I/O on networks of workstations. Parallel and Distributed Computing Practices, 2(2):??, ???? 1999. CO-DEN ???? ISSN 1097-

2803.  URL http://www.
cs.okstate.edu/~pdcp/vols/
vol02/vol02no2abs.html#cheng.

Cong:1999:ENN

[Con99] B. Cong. On encoding neural networks to estimate the atmospheric point spread function in a parallel environment. Parallel and Distributed Computing Practices, 2(4):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol02/vol02no4abs.html#cong.

Cavin:2000:PSL

[CPA00] X. Cavin, J.-C. Paul, and L. Alonso. Partitioning and scheduling large radiosity computations in parallel. Parallel and Distributed Computing Practices, 3(3):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol03/vol03no3abs.html#cavin.

Chatonnay:2000:EMM

[CPC00] P. Chatonnay, L. Philippe, and H. Y. Chan. Evaluation of a multicriteria method to optimize resource access in distributed object systems. Parallel and Distributed Computing Practices, 3(1):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol03/vol03no1abs.html#chatonnay.

Cobb:2000:I

[CS00] Maria Cobb and Kevin Shaw. Introduction. Parallel and Distributed Computing Practices, 3(1):??, ???? 2000. CO-DEN ???? ISSN 1097-

                                        2803.  URL http://www.
                                        cs.okstate.edu/~pdcp/vols/
                                        vol03/vol03no1introduction.
                                        html.

Contassot-Vivier:2002:EDH

[CVLNP02] S. Contassot-Vivier, F. Lombard, J.-M. Nicod, and Laurent Philippe. Evaluation of the DIET hierarchical metacomputing architecture. Parallel and Distributed Computing Practices, 5(4):64–76, December 2002. CODEN ???? ISSN 1097-2803.

Chen:2001:COD

[CX01] Shiping Chen and Jingling Xue. Communication overhead on distributed memory machines. Parallel and Distributed Computing Practices, 4(1):??, ???? 2001. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol04/vol04no1abs.html#chen.

Dupuis:1998:PST

[DC98] Alexandre Dupuis and Bastien Chopard. Parallel simulation of traffic in Geneva using cellular automata. Parallel and Distributed Computing Practices, 1(3):??, ???? 1998. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol01/vol01no3abs.html#dupuis.

DAmbra:2002:PCM

[DdSL02] P. D’Ambra, D. di Serafino, and M. Lapegna. Parallel components for multidimensional quadrature: Some experiences. Parallel and Distributed Computing Practices, 5(3):??, September 2002. CODEN ???? ISSN 1097-2803.

Desprez:2002:ISI

[DFKL02] Frédéric Desprez, Eric Fleury, Alexey Kalinov, and Alexey Lastovetsky. Introduction to the special issue: Algorithms, models and tools for high performance computing on heterogeneous networks. Parallel and Distributed Computing Practices, 5(4):??, December 2002. CODEN ???? ISSN 1097-2803.

Dongarra:1999:P

[DK99] J. Dongarra and E. J. Kontoghiorghes. Preface. Parallel and Distributed Computing Practices, 2(4):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol02/vol02no4preface.html.

deKock:1998:RTS

[dKAE98] E. A. de Kock, Emile H. L.

Aarts, and G. Essink. Realtime scheduling in video systems. Parallel and Distributed Computing Practices, 1(2):??, ???? 1998. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol01/vol01no2abs.html#kock.

Duarte:2001:SLD

[DLC01] Vitor Duarte, João Lourenço, and José C. Cunha. Supporting on-line distributed monitoring and debugging. Parallel and Distributed Computing Practices, 4(3):??, September 2001. CODEN ???? ISSN 1097-2803.

dosSantos:1999:SMS

[dSBA99] R. W. dos Santos, R. Bianchini, and C. L. Amorim. A survey of messaging software issues and systems for Myrinet-based clusters. Parallel and Distributed Computing Practices, 2(2):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol02/vol02no2abs.html#santos.

Kuhn:1998:IHA

[eK98] eva Kühn. Introduction: How to approach the virtual shared memory paradigm. Parallel and Distributed Computing Practices, 1(3):??, ???? 1998. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol01/vol01no3introduction.
                                        html.

Emad:1999:MSD

[Ema99] N. Emad. Mapping strategies in data parallel programming models; the projection methods.

Parallel and Distributed Computing Practices, 2(4):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol02/vol02no4abs.html#emad.

Eberle:2000:SMR

[ES00] H. Eberle and D. Scherrer. A scalable multiprocessor for realtime signal processing applications. Parallel and Distributed Computing Practices, 3(2):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol03/vol03no2abs.html#eberle.

Fischer:1999:EWN

[FD99] M. Fischer and J. Dongarra.

Experiences with Windows NT as a cluster computing platform for parallel computing. Parallel and Distributed Computing Practices, 2(2):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol02/vol02no2abs.html#fido.

Fet:2000:I

[FK00] Kakov Fet and Peter Kacsuk. Introduction. Parallel and Distributed Computing Practices, 3(2):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol03/vol03no2introduction.
html.

Fouad:1999:RTP

[FNH99] H. Fouad, B. Narahari, and J. K. Hahn. A real-time parallel scheduler for the imprecise computation model. Parallel and Distributed Computing Practices, 2(1):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol02/vol02no1abs.html#fouad.

Fry:1999:EDC

[FNS99] P. Fry, J. Nesheiwat, and B. Szymanski. Experiences with distributed computation of twin primes distributions. Parallel and Distributed Computing Practices, 2(3):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol02/vol02no3abs.html#fry.

Fairen:2000:I

[FP00] Marta Fairén and Xavier Pueyo. Introduction. Parallel and Distributed Computing Practices, 3(3):??, ???? 2000. CO-DEN ???? ISSN 1097-

2803.  URL http://www.
cs.okstate.edu/~pdcp/vols/
vol03/vol03no3introduction.
html.

Gannod:1999:RTM

[GB99] B. D. Gannod and S. Bhattacharya. Real-time multicast in wireless communication. Parallel and Distributed Computing Practices, 2(1):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol02/vol02no1abs.html#gannod.

Gustafsson:1998:ADP

[GE98] Jan Gustafsson and Andreas Ermedahl. Automatic derivation of path and loop annotations in object-oriented real- time programs. Parallel and Distributed Computing Practices, 1(2):??, ???? 1998. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol01/vol01no2abs.html#gustafsson.

Gaspar:1999:ADR

[GFS99] C. Gaspar, B. Franek, and J. J. Schwarz. Architecture of a distributed real-time system to control large high-energy physics experiments. Parallel and Distributed Computing Practices, 2(1):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol02/vol02no1abs.html#gaspar.

Gallud:1999:DPR

[GGCM99] J. A. Gallud, J. Garcia-Consuegra, and A. Martinez. Distributed processing of remotely sensed Landsat-TM imagery using MPI. Parallel and Distributed Computing Practices, 2(2):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol02/vol02no2abs.html#gallud.

Gu:2002:BRD

[GH02] Qijun Gu and Danita S. Hill.

Book review: Designing a Wireless Network. Parallel and Distributed Computing Practices, 5 (2):259–260, June 2002. CO-DEN ???? ISSN 1097-2803.

Girau:2000:CCP

[Gir00] B. Girau. Conciliating connectionism and parallel digital hardware. Parallel and Distributed Computing Practices, 3(2):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol03/vol03no2abs.html#girau.

Guyennet:1999:GAC

[GL99] H. Guyennet and J. C. Lapayre. The group approach in cooperative work and in load balancing. Parallel and Distributed Computing Practices, 2(3):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol02/vol02no3abs.html#guyennet.

Gan:2001:DCS

[GLC^{+}01] Boon-Ping Gan, Yoke-Hean Low, Wentong Cai, Stephen J. Turner, Sanjay Jain, Wen Jing Hsu, and Shell Ying Huang. The development of conservative superstep protocols for shared memory multiprocessor systems. Parallel and Distributed Computing Practices, 4(1):??, ???? 2001. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol04/vol04no1abs.html#gan.

Gonzalez:2000:APR

[GMMV00] P. González, E. Moyano, J. P. Molina, and I. Verdú. Analyzing parallel ray tracing algorithms.

Parallel and Distributed Computing Practices, 3(3):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol03/vol03no3abs.html#gonzales.

Gladwell:1998:CPM

[GN98] I. Gladwell and M. A. Napierala. Comparing parallel multidimensional integration algorithms. Parallel and Distributed Computing Practices, 1(1):??, ???? 1998. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol01/vol01no1abs.html#gladwell.

Gomaa:1998:UCD

[Gom98]  Hassan Gomaa.  Use cases

for distributed real-time software architectures. Parallel and Distributed Computing Practices, 1(2):??, ???? 1998. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol01/vol01no2abs.html#gomma.

Goscinski:2001:EPS

[Gos01] Andrzej M. Goscinski. Editorial: A provision of a single system image can move parallel processing on non-dedicated clusters to the mainstream computing. Parallel and Distributed Computing Practices, 4(3):??, September 2001. CODEN ???? ISSN 1097-2803.

Georgiev:1999:PSM

[GZ99] K. Georgiev and Z. Zlatev. Parallel sparse matrix algorithms for air pollution models. Parallel and Distributed Computing Practices, 2(4):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol02/vol02no4abs.html#georgiev.

Hu:1999:LBU

[HB99] Y. F. Hu and R. J. Bake. Load balancing for unstructured mesh applications. Parallel and Distributed Computing Practices, 2(3):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol02/vol02no3abs.html#hu.

Halada:2002:PBQ

[HBL02] L. Halada, S. Benkner, and M. Lucka. Parallel Birge and Qi factorization of three-stage stochastic linear programs. Parallel and Distributed Computing Practices, 5(3):??, September 2002. CODEN ???? ISSN 1097-2803.

He:2002:MOP

[HD02] Yun He and Chris H. Q. Ding. MPI and OpenMP paradigms on cluster of SMP architectures: The vacancy tracking algorithm for multi-dimensional array transposition. Parallel and Distributed Computing Practices, 5(2):117–128, June 2002. CODEN ???? ISSN 1097-2803.

Hyde:1998:CVD

[HF98] Randall L. Hyde and Brett D. Fleish. A case for virtual distributed objects. Parallel and Distributed Computing Practices, 1(3):??, ???? 1998. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol01/vol01no3abs.html#hyde.

Haring:2001:PID

[HHH01] Günter Haring, Helmut Hlavacs, and Karin Hummel. Performance issues in distributed systems. Parallel and Distributed Computing Practices, 4(4):??, December 2001. CODEN ???? ISSN 1097-2803.

Hayder:1999:TPP

[HIK99] M. Ehtesham Hayder, C. S.

Ierotheou, and D. E. Keyes. Three parallel programming paradigms: Comparisons on an archetypal PDE computations. Parallel and Distributed Computing Practices, 2(3):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol02/vol02no3abs.html#hayder.

Hoang:2002:SRT

[HJHK02] H. Hoang, M. Jonsson, U. Hagstorm, and A. Kallerdahl. Switched real-time ethernet with earliest deadline first scheduling— protocols, trafic handling and simuation analysis. Parallel and Distributed Computing Practices, 5(1):105–115, March 2002. CODEN ???? ISSN 1097-2803.

Hawick:2000:RDA

[HJM00] K. A. Hawick, H. A. James, and J. A. Mathew. Remote data access in distributed objectoriented middleware. Parallel and Distributed Computing Practices, 3(1):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol03/vol03no1abs.html#hawick.

Huuck:1998:ITC

[HLL^{+}98] R. Huuck, Y. Lakhnech, B. Lukoschus, L. Urbina, S. Engell, S. Kowalewski, and J. Preuig. Integrating timed condition/event systems and timed automata for the verification of hybrid systems. Parallel and Distributed Computing Practices, 1(2):??, ???? 1998. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol01/vol01no2abs.html#huuck.

Huh:2002:WCA

[HNW^{+}02] E. Huh, A. Neelamegam, Lonnie R. Welch, H. Park, and Barbara Pfarr. Worst-case analysis for real-time processes in time-sharing environments. Parallel and Distributed Computing Practices, 5(1):52–62, March 2002. CODEN ???? ISSN 1097-2803.

Hadjidoukas:2002:MOI

[HPP02] Panagiotis E. Hadjidoukas, Eleftherios D. Polychronopoulos, and Theodore S. Papatheodorou. A modular OpenMP implementation for clusters of multiprocessors. Parallel and Distributed Computing Practices, 5(2):153–168, June 2002. CODEN ???? ISSN 1097-2803.

Hu:2000:OOD

[HPS00] J. C. Hu, I. Pyarali, and D. C.

Schmidt. The object-oriented design and performance of JAWS. Parallel and Distributed Computing Practices, 3(1):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol03/vol03no1abs.html#hu.

Hooman:1998:FDR

[HvR98] Jozef Hooman and Onno van Roosmalen. Formal design of real-time systems in a platformindependent way. Parallel and Distributed Computing Practices, 1(2):??, ???? 1998. CODEN ???? ISSN 1097-

2803.  URL http://www.
cs.okstate.edu/~pdcp/vols/
vol01/vol01no2abs.html#hooman.

Hansen:2002:PCQ

[HZLR02] J. P. Hansen, H. Zhu, J. Lehoczky, and R. Rajkumar. Partitioning considerations for quantized EDF scheduling. Parallel and Distributed Computing Practices, 5(1):63–76, March 2002. CODEN ???? ISSN 1097-2803.

Isabel:2002:CMO

[ICC02] Dorta Isabel, León Coromoto, and Rodríguez Casiano. Comparing MPI and OpenMP implementations of the 0-1 knapsack problem. Parallel and Distributed Computing Practices, 5 (2):129–137, June 2002. CO-DEN ???? ISSN 1097-2803.

Ioannidis:1999:URN

[IPM^{+}99] S. Ioannidis, A. E. Papathanasiou, G. I. Magklis, E. P. Markatos, D. N. Pnevmatikatos, and J. Sevaslidou. On using reliable network RAM in networks of workstations. Parallel and Distributed Computing Practices, 2(2):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol02/vol02no2abs.html#ioan.

Islam:2002:IAC

[ITT02] Mohammad Towhidul Islam, Parimala Thulasiraman, and Ruppa K. Thulasiram. Implementation of ant colony optimization algorithm for mobile ad hoc network applications: OpenMP experiences. Parallel and Distributed Computing Practices, 5(2):177–191, June 2002. CODEN ???? ISSN 1097-2803.

Jiang:2001:ECT

[JVZ01] Shu Jiang, Nitin H. Vaidya, and Wei Zhao. Energy consumption of traffic padding schemes in wireless ad hoc networks. Parallel and Distributed Computing Practices, 4(2):??, ???? 2001. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol04/vol04no2abs.html#jiang.

Jain:2002:MAQ

[JWP^{+}02] Shikha Jain, Lonnie R. Welch, Barbara Pfarr, David M. Chelberg, C. Bruggeman, David Fleeman, David Parrott, Zhenyu Tan, M. McCoy, and Chris Shuler. Management of application QOS-level adaptation for real-time systems. Parallel and Distributed Computing Practices, 5(1):15–26, March 2002. CODEN ???? ISSN 1097-2803.

Karatza:2001:GSP

[Kar01] Helen Karatza. Gang scheduling performance under different distributions of gang size. Parallel and Distributed Computing Practices, 4(4):??, December 2001. CODEN ???? ISSN 1097-2803.

Kose:2000:CME

[KC00] C. Köse and A. Chalmers. Clusteral models for efficient parallel volume visualization. Parallel and Distributed Computing Practices, 3(3):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol03/vol03no3abs.html#kose.

Keyes:1999:EAH

[Key99] David Keyes. Editorial: Adapting to hostile architectural environments. Parallel and Distributed Computing Practices, 2(1):??, ???? 1999. CO-DEN ???? ISSN 1097-

2803.  URL http://www.
cs.okstate.edu/~pdcp/vols/
vol02/vol02no1editorial.html.

Kimm:2001:BRP

[Kim01] Haklin Kimm. Book review:

Parallel Computing Using Optical Interconnections. Parallel and Distributed Computing Practices, 4(4):??, December 2001. CODEN ???? ISSN 1097-2803.

Kacsuk:2001:ISI

[KK01a] Péter Kacsuk and Gabriele Kotsis. Introduction to the special issue: Quality of parallel and distributed programs and systems. Parallel and Distributed Computing Practices, 4(4):??, December 2001. CODEN ???? ISSN 1097-2803.

Kovacs:2001:DDD

[KK01b] Jozsef Kovács and Péter Kacsuk. The DIWIDE distributed debugger. Parallel and Distributed Computing Practices, 4 (4):??, December 2001. CODEN ???? ISSN 1097-2803.

Kowalik:1999:EWC

[KL99] J. Kowalik and J. Lixvar. Editorial: A wakeup call for supercomputing. Parallel and Distributed Computing Practices, 2(4):??, ???? 1999. CODEN ???? ISSN 1097-

                                        2803.  URL http://www.
                                        cs.okstate.edu/~pdcp/vols/
                                        vol02/vol02no4editorial.html.

Karavanic:2001:FME

[KM01] Karen L. Karavanic and Barton P. Miller. A framework for multi-execution performance tuning. Parallel and Distributed Computing Practices, 4(3):??, September 2001. CODEN ???? ISSN 1097-2803.

Ko:2002:CBA

[KM02] Ren-Song Ko and Matt W.

Mutka. A component-based approach for adaptive soft realtime Java within heterogeneous environments. Parallel and Distributed Computing Practices, 5 (1):89–104, March 2002. CO-DEN ???? ISSN 1097-2803.

Kopyssov:2002:PAM

[KN02] Sergey P. Kopyssov and Alexander K. Novikov. Parallel adaptive mesh refinement with load balancing on heterogeneous cluster. Parallel and Distributed Computing Practices, 5(4):??, December 2002. CODEN ???? ISSN 1097-2803.

Kotov:2000:FAO

[Kot00] Vadim Kotov. Facilitating analysis and optimization of largescale enterprise systems. Parallel and Distributed Computing Practices, 3(4):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol03/vol03no4abs.html#kotov.

Kowalik:2000:SIE

[Kow00] Janusz Kowalik. Special issue Editor introduction. Parallel and Distributed Computing Practices, 3(4):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol03/vol03no4introduction.
html.

Kewley:2001:FFI

[KP01] John M. Kewley and Radu Prodan. The FIRST framework for interoperable run-time software tools. Parallel and Distributed Computing Practices, 4 (3):??, September 2001. CO-DEN ???? ISSN 1097-2803.

Korosec:2002:ESA

[KSR02] P. Korosec, J. Silc, and B. Robic. An experimental study of an ant-colony algorithm for the mesh-partitioning problem. Parallel and Distributed Computing Practices, 5(3):??, September 2002. CODEN ???? ISSN 1097-2803.

Konstantinopulos:2000:EIC

[KSSM00] C. Konstantinopulos, A. Svolos, D. Serpanos, and D. Maritsas. The effect of interword connectivity in associative processing. Parallel and Distributed Computing Practices, 3(2):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol03/vol03no2abs.html#konstan.

Kutil:2002:OBA

[Kut02] R. Kutil. Optimization of bitstream assembly in parallel multimedia compression. Parallel and Distributed Computing Practices, 5(3):??, September 2002. CODEN ???? ISSN 1097-2803.

Kostkova:1998:VST

[KW98] Patty Kostkova and Tim Wilkinson. A virtual shared tuplespace resource manager. Parallel and Distributed Computing Practices, 1(3):??, ???? 1998. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol01/vol01no3abs.html#kostkova.

Lopez:2002:ESM

[LA02] Félix César García López and Nieves Luz Frías Arrocha. Expanding the synchronization model for OpenMP. Parallel and Distributed Computing Practices, 5(2):169–175, June 2002. CODEN ???? ISSN 1097-2803.

Lin:1999:TSB

[LG99] W-. M. Lin and Q. Gu. Task scheduling on bus-based networks of workstations. Parallel and Distributed Computing Practices, 2(2):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol02/vol02no2abs.html#lin.

Lim:2001:EIS

[LLS01] Sung-Soo Lim, Kihwal Lee, and Lui Sha. Ensuring integrity and service availability in a Web-based control laboratory. Parallel and Distributed Computing Practices, 4(2):??, ???? 2001. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol04/vol04no2abs.html#lim.

Ludwig:2001:ISI

[LM01] Thomas Ludwig and Barton P.

Miller. Introduction to the special issue: Monitoring systems and tool interoperability. Parallel and Distributed Computing Practices, 4(3):??, September 2001. CODEN ???? ISSN 1097-2803.

Liang:1999:MLW

[LS99] W. Liang and H. Shen. Multicast in large WDM networks. Parallel and Distributed Computing Practices, 2(3):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol02/vol02no3abs.html#liang.

Lin:2002:LSH

[LS02] Wei-Ming Lin and Yongsheng Song. Loop scheduling for heterogeneous computing environment based on performance prediction. Parallel and Distributed Computing Practices, 5(2):219– 231, June 2002. CODEN ???? ISSN 1097-2803.

Lumpp:1998:PVD

[LSDG98] James E. Lumpp, Jr., Kuppuswamy Sivakumar, Christopher Diaz, and James N. Griffioen. Performance visualization for distributed shared memory systems. Parallel and Distributed Computing Practices, 1(3):??, ???? 1998. CO-DEN ???? ISSN 1097-

                                        2803.  URL http://www.
                                        cs.okstate.edu/~pdcp/vols/
                                        vol01/vol01no3abs.html#lumpp.

Li:2001:DTC

[LY01] Xiangyang Li and Nong Ye. Decision tree classifiers for computer intrusion detection. Parallel and Distributed Computing Practices, 4(2):??, ???? 2001. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol04/vol04no2abs.html#li.

MacIntyre:2000:EPD

[Mac00] B. MacIntyre. Exploratory programming of distributed 3D graphics applications. Parallel and Distributed Computing Practices, 3(3):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol03/vol03no3abs.html#macintyre.

Marowka:2002:EWG

[Mar02a] Ami Marowka. Editorial: What is the GRID? Parallel and Distributed Computing Practices, 5 (1):iii–iv, March 2002. CODEN ???? ISSN 1097-2803.

Marowka:2002:ISI

[Mar02b] Ami Marowka. Introduction to the special issue: OpenMP: Experiences, implementations and applications. Parallel and Distributed Computing Practices, 5 (2):v, June 2002. CODEN ???? ISSN 1097-2803.

Masters:2002:ENC

[Mas02] Michael W. Masters. Editorial: New challenges in managing distributed systems. Parallel and Distributed Computing Practices, 5(3):??, September 2002. CODEN ???? ISSN 1097-2803.

Moradi:2001:SOM

[MAT01] Farshad Moradi, Rassul Ayani, and Gary Tan. Some ownership management issues in distributed simulation using HLA/ RTI. Parallel and Distributed Computing Practices, 4(1):??, ???? 2001. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol04/vol04no1abs.html#moradi.

Marowka:2000:SPM

[MB00] A. Marowka and M. Bercovier. A scalable portability model for parallel computing. Parallel and Distributed Computing Practices, 3(3):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol03/vol03no3abs.html#marowka.

Milutinovic:1998:EST

[Mil98]  Veljko Milutinovic.  Edito-

rial: Some thoughts about writing of journal papers. Parallel and Distributed Computing Practices, 1(3):??, ???? 1998. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol01/vol01no3editorial.html.

Miyoshi:1998:RTJ

[MKT98] Akihiko Miyoshi, Takuro Kitayama, and Hideyuki Tokuda. A real-time Java server for realtime Mach. Parallel and Distributed Computing Practices, 1(2):??, ???? 1998. CO-DEN ???? ISSN 1097-

2803.  URL http://www.

cs.okstate.edu/~pdcp/vols/
vol01/vol01no2abs.html#miyoshi.

Maier:1999:ESR

[MLB99] U. Maier, P. Luksch, and A. Bode. Experiences with the SEMPA resource manager for real world scientific computing in networks of workstations. Parallel and Distributed Computing Practices, 2(2):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol02/vol02no2abs.html#maier.

Manimaran:1998:IST

[MMMM98] G. Manimaran, Shashidhar Merugu, Anand Manikutty, and C. Siva Ram Murthy. Integrated scheduling of tasks and messages in distributed real-time systems. Parallel and Distributed Computing Practices, 1(2):??, ???? 1998. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol01/vol01no2abs.html#manimaran.

Marcoux:1998:GOD

[MMMS98] Annie Marcoux, Christine Maurel, Frederic Migeon, and Patrick Salle. Generic operational decomposition for concurrent systems: Semantics and reflection. Parallel and Distributed Computing Practices, 1(4):??, ???? 1998. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol01/vol01no4abs.html#migeon.

Mahjoub:2002:CIT

[MNT02] Zaher Mahjoub, Wahid Nasri, and Denis Trystram. Computing the inverse of a triangular matrix on heterogeneous clusters. Parallel and Distributed Computing Practices, 5(4):??, December 2002. CODEN ???? ISSN 1097-2803.

Mancini:2000:LHP

[MP00] S. Mancini and R. Pacalet. LUX: A heterogeneous parallel computer dedicated to ray tracing. Parallel and Distributed Computing Practices, 3(3):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol03/vol03no3abs.html#mancini.

Malony:2001:PTC

[MS01] Allen D. Malony and Sameer Shende. Performance technology for complex parallel and distributed systems. Parallel and Distributed Computing Practices, 4(4):??, December 2001. CODEN ???? ISSN 1097-2803.

Maniatty:1998:PCG

[MSC98] W. A. Maniatty, B. K. Szymanski, and T. Caraco. Parallel computing with generalized cellular automata. Parallel and Distributed Computing Practices, 1(1):??, ???? 1998. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol01/vol01no1abs.html#maniatty.

Molla:2000:PFP

[MV00] R. Mollá and R. Vivó. Parallel fixed-point digital differential analyser with antialiasing PFD-DAA. Parallel and Distributed Computing Practices, 3(3):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol03/vol03no3abs.html#molla.

Masters:2001:ECB

[MW01] Michael W. Masters and Lonnie R. Welch. Editorial: Challenges for building complex realtime computing systems. Parallel and Distributed Computing Practices, 4(2):??, ???? 2001. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol04/vol04no2editorial.html.

Nagasawa:2000:WIS

[Nag00] Mikio Nagasawa. What-if scenario explorer for performance and scalability simulations. Parallel and Distributed Computing Practices, 3(4):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol03/vol03no4abs.html#nagasawa.

Naim:1998:VDL

[NH98] O. Naim and A. J. G. Hey.

Visualization of do-loop performance using AVS. Parallel and Distributed Computing Practices, 1(1):??, ???? 1998. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol01/vol01no1abs.html#naim.

Nezlek:2000:ESC

[NJN00] George S. Nezlek, Hemant K. Jain, and Derek. L. Nazareth. Estimating server capacities for enterprise computing architectures. Parallel and Distributed Computing Practices, 3(4):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol03/vol03no4abs.html#nezlek.

Ng:1998:EDC

[NS98] Joseph Kee-Yin Ng and Shibin Song. Efficient delay computation methods for real-time traffic in an ATM network. Parallel and Distributed Computing Practices, 1(4):??, ???? 1998. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol01/vol01no4abs.html#ng.

Ng:1998:DFT

[NsL98] Joseph Kee-Yin Ng and Victor Chung sing Lee. Delayed frame transmission schemes for MPEG videos in a real-time. Parallel and Distributed Computing Practices, 1(3):??, ???? 1998. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol01/vol01no3abs.html#joseph.

Ng:1999:DIV

[NX99] Joseph Kee-Ying Ng and Shuhua Xiong. The design and implementation of video servers in video-on-demand systems. Parallel and Distributed Computing Practices, 2(3):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol02/vol02no3soft_reviews.
html#ng   target=.

Oksa:2002:PCS

[OBV02] Gabriel Oksa, Martin Becka, and Marián Vajtersic. Parallel computation with structured matrices in linear modeling of multidimensional signals. Parallel and Distributed Computing Practices, 5(3):??, September 2002. CODEN ???? ISSN 1097-2803.

Omma:2001:BRS

[Omm01] Mikael Omma. Book review:

Server Based Java Programming. Parallel and Distributed Computing Practices, 4(4):??, December 2001. CODEN ???? ISSN 1097-2803.

Olivares:2002:PSN

[OOBG^{+}02] T. Olivares, L. Orozco-Barbosa, P. J. Garcia, F. Quiles, and A. Garrido. On the performance and scalability of NFS over a Myrinet-based network of workstations. Parallel and Distributed Computing Practices, 5 (2):233–238, June 2002. CO-DEN ???? ISSN 1097-2803.

Overeinder:2001:SOC

[OSS01] B. J. Overeinder, A. Schoneveld, and P. M. A. Sloot. Self-organized criticality in optimistic simulation of correlated systems. Parallel and Distributed Computing Practices, 4(1):??, ???? 2001. CODEN ???? ISSN 1097-

                                        2803.  URL http://www.
                                        cs.okstate.edu/~pdcp/vols/
                                        vol04/vol04no1abs.html#overeinder.

Pal:2001:BRP

[Pal01] Michelle Pal. Book review: Parallel Database Techniques. Parallel and Distributed Computing Practices, 4(3):??, September 2001. CODEN ???? ISSN 1097-2803.

Pan:2000:BRa

[Pan00a] Yi Pan. Book reviews. Parallel and Distributed Computing Practices, 3(3):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol03/vol03no3reviews.html.

Pan:2000:BRb

[Pan00b] Yi Pan. Book reviews. Parallel and Distributed Computing Practices, 3(4):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol03/vol03no4reviews.html.

Pan:2001:BR

[Pan01] Yi Pan. Book reviews. Parallel and Distributed Computing Practices, 4(2):??, ???? 2001. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol04/vol04no2reviews.html.

Paprzycki:2000:RV

[Pap00]  Marcin Paprzycki.  Refer-

ees for volume 3. Parallel and Distributed Computing Practices, 3(4):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol03/vol03no4referees.html.

Paprzycki:2001:EWN

[Pap01] Marcin Paprzycki. Editorial:

What’s in a name. Parallel and Distributed Computing Practices, 4(4):??, December 2001. CODEN ???? ISSN 1097-2803.

Podhorszki:2001:SLM

[PK01] Norbert Podhorszki and Péter Kacsuk. Semi-on-line monitoring of P-GRADE applications. Parallel and Distributed Computing Practices, 4(4):??, December 2001. CODEN ???? ISSN 1097-2803.

Periquet:2000:DPM

[PL00] A. I. Periquet and E. C. Lin.

The design and performance of a meta-level mobile component for video delivery. Parallel and Distributed Computing Practices, 3(1):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol03/vol03no1abs.html#periquet.

Patnaik:1998:IFP

[PMM98] L. M. Patnaik, Ravi Mittal, and K. M. Mehta. Introduction: Facilitating parallel and distributed computing. Parallel and Distributed Computing Practices, 1(4):??, ???? 1998. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol01/vol01no4introduction.
html.

Pschernig:2002:PAO

[PU02] E. Pschernig and Andreas Uhl. Parallel algorithms for overlapped block-matching motion compensation. Parallel and Distributed Computing Practices, 5 (3):??, September 2002. CO-DEN ???? ISSN 1097-2803.

Polze:1999:RTC

[PWPM99] A. Polze, K. Wallnau, D. Plakosh, and M. Malek. Real-time computing with off-the-shelf components: The case for CORBA. Parallel and Distributed Computing Practices, 2(1):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol02/vol02no1abs.html#polze.

Pfarr:2002:ISI

[PZB02] Barbara Pfarr, Armin Zimmerman, and Scott Brandt. Introduction to the special issue: Parallel and distributed realtime systems. Parallel and Distributed Computing Practices, 5 (1):v–vi, March 2002. CODEN ???? ISSN 1097-2803.

Qureshi:2000:EST

[QH00] K. Qureshi and M. Hatanaka. An empirical study of task scheduling strategies for image processing application on heterogeneous distributed computing system. Parallel and Distributed Computing Practices, 3(3):??, ???? 2000. CO-DEN ???? ISSN 1097-

2803.  URL http://www.
cs.okstate.edu/~pdcp/vols/
vol03/vol03no3abs.html#qureshi.

Reinhard:2000:SDI

[RCJ00] E. Reinhard, A. Chalmers, and F. W. Jansen. Sampling diffuse inter-reflection within a hybrid scheduling ray tracer. Parallel and Distributed Computing Practices, 3(3):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol03/vol03no3abs.html#reinhard.

Reddaway:1999:EWS

[Red99] Stewart Reddaway. Editorial:

Will SIMD make a comeback? Parallel and Distributed Computing Practices, 2(3):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol02/vol02no3editorial.html.

Ribeiro-Justo:1998:AFD

[RJC98] G. R. Ribeiro-Justo and P. R. F.

Cunha. An application framework for dynamic distributed software architectures. Parallel and Distributed Computing Practices, 1(4):??, ???? 1998. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol01/vol01no4abs.html#justo.

Rukoz:2000:SJT

[RLR00] M. Rukoz, C. Leon, and M. Rivas. SIMA: A Java tool for constructing image processing applications on a heterogeneous network. Parallel and Distributed Computing Practices, 3(1):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol03/vol03no1abs.html#rukoz.

Raje:1998:DIC

[RMB^{+}98] R. R. Raje, S. Mukhopadhyay, M. Boyles, N. Patel, and A. Papiez. On designing and implementing a collaborative system using the distributed-object model of Java-RMI. Parallel and Distributed Computing Practices, 1(4):??, ???? 1998. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol01/vol01no4abs.html#raje.

Sajeev:1998:PPU

[Saj98] A. S. M. Sajeev. Parallel programming using visual patterns. Parallel and Distributed Computing Practices, 1(4):??, ???? 1998. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol01/vol01no4abs.html#sajeev.

Sandhu:1999:MWE

[SBM99] H. Sandhu, T. Brecht, and D. Moscoso. Multiple-writer entry consistency. Parallel and Distributed Computing Practices, 2(2):??, ???? 1999.

CODEN ???? ISSN 1097-

2803.  URL http://www.
cs.okstate.edu/~pdcp/vols/
vol02/vol02no2abs.html#sand.

Schmidt:2000:ETD

[Sch00]  Douglas C. Schmidt.  Edi-

torial: Trends in distributed object computing. Parallel and Distributed Computing Practices, 3(1):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol03/vol03no1editorial.html.

Schurer:2002:OCI

[Sch02]  R. Schurer.  Optimal com-

munication interval for parallel adaptive integration. Parallel and Distributed Computing Practices, 5(3):??, September 2002. CODEN ???? ISSN 1097-2803.

Snell:1999:DAP

[SJC99] Q. Snell, G. Judd, and M. Clement.

The DOGMA approach to parallel and distributed computing. Parallel and Distributed Computing Practices, 2(2):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol02/vol02no2abs.html#snell.

Shi:2001:ROF

[SL01] Hongchi Shi and Hongzheng Li. Rank order filtering on massively-parallel single-bit mesh processor arrays. Parallel and Distributed Computing Practices, 4(1):??, ???? 2001. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol04/vol04no1abs.html#shi.

Swaminathan:2002:FFB

[SM02a] S. Swaminathan and G. Manimaran. FARM: A feedbackbased adaptive resource management for complex real-time systems and application to sensor Web. Parallel and Distributed Computing Practices, 5 (1):1–13, March 2002. CODEN ???? ISSN 1097-2803.

Swaminathan:2002:VBS

[SM02b] S. Swaminathan and G. Manimaran. A value-based scheduler capturing schedulabilityreliability tradeoff in multiprocessor real-time systems. Parallel and Distributed Computing Practices, 5(1):77–87, March 2002. CODEN ???? ISSN 1097-2803.

Song:2000:EDE

[Son00] S. W. Song. Editorial: Design of efficient and scalable parallel algorithms. Parallel and Distributed Computing Practices, 3(3):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol03/vol03no3editorial.html.

Subramanian:1998:FPB

[SP98] Ram Subramanian and Santosh Pande. A framework for performance-based program partitioning. Parallel and Distributed Computing Practices, 1(4):??, ???? 1998. CO-DEN ???? ISSN 1097-

2803.  URL http://www.
cs.okstate.edu/~pdcp/vols/
vol01/vol01no4abs.html#ram.

Silc:1998:APC

[SRU98] J. Silc, B. Robic, and T. Ungerer. Asynchrony in parallel computing: From dataflow to multithreading. Parallel and Distributed Computing Practices, 1(1):??, ???? 1998. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol01/vol01no1abs.html#silc.

Subramanian:2001:AMP

[SRW01] S. Subramanian, D. M. Rao, and P. A. Wilsey. Applying multilevel partitioning to parallel logic simulation. Parallel and Distributed Computing Practices, 4(1):??, ???? 2001. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol04/vol04no1abs.html#subramanian.

Schmidt:2000:KIS

[SS00] B. Schmidt and M. Schimmler. KPROC — an instruction systolic architecture for parallel prefix applications. Parallel and Distributed Computing Practices, 3(2):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol03/vol03no2abs.html#schmidt.

Spezzano:1998:CPE

[ST98] Giandomenico Spezzano and Domenico Talia. The CAR-PET programming environment for solving scientific problems on parallel computers. Parallel and Distributed Computing Practices, 1(3):??, ???? 1998. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol01/vol01no3abs.html#spezzano.

Sterk:2002:NSF

[STP02] M. Sterk, Roman Trobec, and M. Praprotnik. Numerical schemes for fluid flow and heat transfer in medical simulations. Parallel and Distributed Computing Practices, 5(3):??, September 2002. CODEN ???? ISSN 1097-2803.

Smith:2000:PSD

[SW00] Connie U. Smith and Lloyd G. Williams. Performance and scalability of distributed software architectures: An SPE approach. Parallel and Distributed Computing Practices, 3(4):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol03/vol03no4abs.html#smith.

Shirazi:2000:DBS

[SWR^{+}00] B. Shirazi, L. Welch, B. Ravindran, C. Cavanaugh, and E. Huh. DynBench: A benchmark suite for dynamic realtime systems. Parallel and Distributed Computing Practices, 3(1):??, ???? 2000.

CODEN ???? ISSN 1097-

2803.  URL http://www.
cs.okstate.edu/~pdcp/vols/
vol03/vol03no1abs.html#shirazi.

Szymanski:1998:ECP

[Szy98] Boleslaw K. Szymanski. Editorial: Challenges in parallel and distributed computing. Parallel and Distributed Computing Practices, 1(4):??, ???? 1998. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol01/vol01no4editorial.html.

Tunali:1999:HFT

[TES99] T. Tunali, K. Erciye, and Z. Soysert. A hierarchical faulttolerant ring protocol for distributed real-time systems. Parallel and Distributed Computing Practices, 2(1):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol02/vol02no1abs.html#tunali.

Tadonki:1999:PMV

[TP99] C. Tadonki and B. Philippe.

Parallel multiplication of a vector by a Kronecker product of matrices. Parallel and Distributed Computing Practices, 2(4):??, ???? 1999. CO-DEN ???? ISSN 1097-

2803.  URL http://www.
cs.okstate.edu/~pdcp/vols/
vol02/vol02no4abs.html#tadonki.

Tadonki:2000:PMV

[TP00] C. Tadonki and B. Philippe.

Parallel multiplication of a vector by a Kronecker product of matrices (Part II). Parallel and Distributed Computing Practices, 3(3):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol03/vol03no3abs.html#tadonki.

Tropper:2001:SIE

[Tro01] Carl Tropper. Special issue Editor introduction. Parallel and Distributed Computing Practices, 4(1):??, ???? 2001. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol04/vol04no1introduction.
                                        html.

Trobec:2002:EPN

[Tro02] Roman Trobec. Editorial: Parallel numeric algorithms on faster computers. Parallel and Distributed Computing Practices, 5 (2):iii–iv, June 2002. CODEN ???? ISSN 1097-2803.

Trapp:2002:ABE

[TST02] M. Trapp, B. Schuermann, and T. Tetteroo. Analyzing the behavior of embedded systems concerning gracefull degradation. Parallel and Distributed Computing Practices, 5(1):27– 40, March 2002. CODEN ???? ISSN 1097-2803.

Thulasiraman:2002:EMA

[TTKG02] Parimala Thulasiraman, Kevin Theobald, Ashfaq A. Khokhar, and Guang R. Gao. Efficent multithreaded algorithms for the Fast Fourier Transform. Parallel and Distributed Computing Practices, 5(2):239–258, June 2002. CODEN ???? ISSN 1097-2803.

Tanev:2000:DBP

[TUO00] I. T. Tanev, T. Uaozumi, and K. Ono. DCOM-based parallel distributed implementation of GP. Parallel and Distributed Computing Practices, 3(1):??, ???? 2000. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol03/vol03no1abs.html#tanev.

Tjaden:2001:SMC

[TW01] Brett Tjaden and Lonnie R.

Welch. Security for mission critical real-time systems. Parallel and Distributed Computing Practices, 4(2):??, ???? 2001. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol04/vol04no2introduction.
html.

Trobec:2002:ISI

[TZVU02] Roman Trobec, Peter Zinterhof, Marián Vajtersic, and Andreas Uhl. Introduction to the special issue: Parallel numerics and applications. Parallel and Distributed Computing Practices, 5 (3):??, September 2002. CO-DEN ???? ISSN 1097-2803.

Vardanega:1999:DCF

[Var99] T. Vardanega. On the distribution of control functions in new-generation on-board embedded real-time systems. Parallel and Distributed Computing Practices, 2(1):??, ????

1999. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol02/vol02no1abs.html#varda.

vanKatwijk:1999:SDV

[vKdRST99] J. van Katwijk, R. de Rooij, S. Stuurman, and H. Toetenel. Software development and verification of dynamic real-time distributed systems based on the radio broadcast paradigm. Parallel and Distributed Computing Practices, 2(1):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol02/vol02no1abs.html#katwijk.

vanKatwijk:1999:PDR

[vKZ99] Jan van Katwijk and Janusz Zalewski. Parallel and distributed real-time systems: An introduction. Parallel and Distributed Computing Practices, 2(1):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol02/vol02no1introduction.
                                        html.

Wu:1999:JBD

[WCS99] X. Wu, Q. Chen, and X.-H. Sun. A Java-based distributed debbuger supporting MPI and PVM. Parallel and Distributed Computing Practices, 2(4):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol02/vol02no4abs.html#wu.

Welch:1998:EOP

[Wel98] Lonnie R. Welch. Editorial:

Open problems in the distributed control systems domain. Parallel and Distributed Computing Practices, 1(2):??, ???? 1998. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol01/vol01no2editorial.html.

Weissman:1999:HPT

[WG99] B. Weissman and B. Gomes.

High performance thread migration on clusters of SMPs. Parallel and Distributed Computing Practices, 2(2):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol02/vol02no2abs.html#boris.

Welch:1998:IED

[WH98] Lonnie R. Welch and Dieter K.

Hammer. Introduction: Engineering of distributed control systems. Parallel and Distributed Computing Practices, 1(2):??, ???? 1998. CO-DEN ???? ISSN 1097-

2803.  URL http://www.
cs.okstate.edu/~pdcp/vols/
vol01/vol01no2introduction.
html.

Weimin:1998:NFM

[WJJ98] Zheng Weimin, Shen Jun, and Wang Ju. A new fast message passing communication system for multiprocessor workstation clusters. Parallel and Distributed Computing Practices, 1(4):??, ???? 1998. CODEN ???? ISSN 1097-2803. URL http://www.

                                        cs.okstate.edu/~pdcp/vols/
                                        vol01/vol01no4abs.html#zhen.

Wu:2002:ALB

[WJT02] Song Wu, Hai Jin, and Guang Tan. Analysis of load balancing issues caused by intramovie skewness for parallel video servers. Parallel and Distributed Computing Practices, 5(2):205– 217, June 2002. CODEN ???? ISSN 1097-2803.

Wismuller:2001:MCP

[WLKB01] Roland Wismuller, Thomas Ludwig, Wolfgang Karl, and Arndt Bode. Monitoring concepts for parallel systems—an evolution towards interoperable tool environments. Parallel and Distributed Computing Practices, 4(3):??, September 2001. CODEN ???? ISSN 1097-2803.

Wolf:2001:SPP

[WM01] Felix Wolf and Bernd Mohr. Specifying performance properties of parallel applications using compound events. Parallel and Distributed Computing Practices, 4(3):??, September 2001. CODEN ???? ISSN 1097-2803.

Wojcik:1999:DCR

[WZ99] Z. M. Wójcik and J. Zalewski. Distributed computations in real time based on a rough grammar principle. Parallel and Distributed Computing Practices, 2(1):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol02/vol02no1abs.html#wojcik.

Yamane:1998:AEV

[Yam98] Satoshi Yamane. Automatic effective verification method for distributed and concurrent systems using timed language inclusion. Parallel and Distributed Computing Practices, 1(2):??, ???? 1998. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol01/vol01no2abs.html#yamane.

Zalewski:2000:ETS

[Zal00] J. Zalewski. Editorial: Tool support for distributed realtime software. Parallel and Distributed Computing Practices, 3(2):??, ???? 2000. CODEN ???? ISSN 1097-

2803.  URL http://www.
cs.okstate.edu/~pdcp/vols/
vol03/vol03no2editorial.html.

Zinterhof:2002:NIH

[Zin02] Peter Zinterhof. Numerical integration and Hilbert–Schmidt norms. Parallel and Distributed Computing Practices, 5(3):??, September 2002. CODEN ???? ISSN 1097-2803.

Zoltani:2001:EPO

[ZSnH01] Csaba K. Zoltani, Punyam Satya-narayana, and Dixie Hisley. Evaluating performance of OpenMP and MPI on the SGI Origin 2000 with benchmarks of realistic problem sizes. Parallel and Distributed Computing Practices, 4(4):??, December 2001. CODEN ???? ISSN 1097-2803.